# A lower bound for classical Kloosterman sums and an application

### v3

@@ Abstract.
We present a lower bound for the classical Kloosterman sum $S(a,b;c)$ where $(ab,c)=1$ and $c$ is an odd integer. We apply this lower bound for Kloosterman sums to derive an explicit lower bound in Petersson’s trace formula, subject to a given condition. Consequently, we achieve a modified version of [JS20, Theorem 1.7], where weight $k$ and level $N$ are permitted to vary independently. Using this modified version, we get a lower bound concerning the trace of the Hecke operator $T_{n}$ acting on the space $S(N,k)$, of cusp forms of weight $k$ and level $N$ for $(n,N)=1$.
@@ 1. Introduction
where $a,b$ are integers, $c$ is a natural number and $\overline{x}(c)$ denotes the multiplicative inverse of $x$ modulo $c$. The sums for the special case of $a=0$ or $b=0$ are called Ramanujan sums. Kloosterman sums appear in various problems in number theory. One of the important places of its occurrence is in the Kuznetsov trace formula [Kuz80, Theorem 1] and Petersson’s trace formula. Recently, the non-vanishing aspect of Kloosterman sums has been used in [BDS23, Theorem 2] and [Das24] to obtain discrepancy results for the distribution of eigenvalues of Hecke operators. Whenever we have non-vanishing of a sum, a natural question to ask is whether we can obtain a lower bound for the same. It appears that there is no such lower bound for Kloosterman sums readily available in the literature. In the next section, we answer this question for $S(a,b;c),$ where $(ab,c)=1$. First, we obtain the following lower bound for classical Kloosterman sums.
@@
Let $c$ be an odd natural number such that $c=df$ where $f$ is squarefree and $d$ is a powerful number with $(d,f)=1$. Let $a,b$ be integers such that $(ab,c)=1$ and $ab$ is a quadratic residue modulo each prime dividing $d$ (and hence, $ab$ is a square modulo $d$ by Hensel’s lemma and Chinese remainder theorem). Then
$\left|S(a,b;c)\right|\geq\frac{2^{\omega(d)}}{\sqrt{d}}\left(\frac{1}{\left(\tau(f)\sqrt{f}\right)^{\phi(f)/\tau(f)-1}}\right),$
where $\omega(d)$ is the number of distinct prime divisors of $d$, $\tau(f)$ is the number of positive divisors of $f,$ and $\phi(f)$ denotes the Euler’s phi function of $f$. If $ab$ is a quadratic non-residue modulo a prime dividing $d$ (and hence $ab$ is not a square modulo $d$), then $S(a,b;c)=0$.
@@
Let $m,n,N$ be natural numbers with $(mn,N)=1$ and $N$ odd. Let $H_{0}=1-\frac{4}{9}-\log\left(\frac{9}{5}\right)\approx-0.03223$, $D_{0}\in\left(e^{H_{0}},1\right)$ be a real number ($e^{H_{0}}\approx 0.96828)$. Let $N=df$ where $d$ is powerful and $f$ is squarefree, where $(d,f)=1$. Let $k_{0}(N)=2+\left\lfloor\frac{\log 7-\log H(N)}{\log A_{0}}\right\rfloor,$ where $A_{0}=D_{0}e^{-H_{0}},$ $H(N)=2^{\omega(N)}\frac{G(N)}{N^{3/2}}$ with $G(N)=\frac{f}{(\tau(f)^{2}f)^{\phi(f)/2\tau(f)}}.$ Let $k\geq\max(22,k_{0}(N))$, $\frac{4\pi\sqrt{mn}}{N}\in\left[D_{0}(k-1),(k-1)\right]$ and assume that $mn$ is a quadratic residue modulo each prime dividing $d$. Let $\Delta_{k,N}(m,n)$ be as defined in equation (3.1), then
@@
  • The theorems [JS20, Theorem 1.7] and [Das24, Theorem 4], assume $N$ to be fixed and provide an asymptotic for $|\Delta_{k,N}(m,n)-\delta(m,n)|,$ as $k\rightarrow\infty.$ However, Theorem 2 is applicable for every $k$ and $N$, provided they satisfy certain hypotheses. For a fixed level $N,$ [ILS00, Corollary 2.3] gives an asymptotic for $\Delta_{k,N}(m,n)$ with the condition $\frac{4\pi\sqrt{mn}}{N}\leq\frac{k}{3},$ whereas Theorem 2 gives us an asymptotic for the condititon $D_{0}(k-1)\leq\frac{4\pi\sqrt{mn}}{N}\leq(k-1).$
@@
The following corollaries are immediate.
@@
Let $0.44730<\epsilon<$ 0.4473 . Let $m,n,N$ be natural numbers with $(mn,N)=1$ and $N$ odd. Let $H_{0},D_{0},k_{0}(N)$ be as defined in Theorem 2. Let $N=df$ where $d$ is powerful and $f$ is squarefree where $(d,f)=1$. If $\frac{4\pi\sqrt{mn}}{N}\in\left[D_{0}(k-1),(k-1)\right]$ and $mn$ is a quadratic residue modulo each prime dividing $d$, then for all $k\geq\max\left(k_{0}(N),1+\left(\frac{3}{\epsilon}\right)^{3/2}\right),$ we have
@@
Let $\mathcal{F}_{k}(N)$ be an orthonormal basis of $S_{k}(N)$ consisting only of joint eigenfunctions of the Hecke operators $T_{n}$. Let $\lambda_{n}(f)$ be the $n$-th normalised Hecke eigenvalue of a cusp form $f.$
Corollary 2.
Let $k,\,N$ be natural numbers with $k\geq\max(22,k_{0}(N)),$ $k$ even and $N$ odd. Let $n\geq 2$ be a natural number such that $n\in\left[\left(\frac{0.999(k-1)N}{4\pi}\right)^{2},\left(\frac{(k-1)N}{4\pi}\right)^{2}\right]$ with $(n,N)=1$, and $n$ is a quadratic residue modulo each prime dividing the powerful part of $N$. Then for $t>1,$ we have
$\sum_{f\in\mathcal{F}_{k}(N)}|\lambda^{t}_{n}(f)|\geq\left(\frac{(4\pi)^{k-1}}{\Gamma(k-1)}\right)^{t}\frac{1}{|\mathcal{F}_{k}(N)|^{t-1}}\left|\frac{0.002397\pi}{(k-1)^{1/3}}e^{(k-1)\left(1-\frac{4}{9}-\log(\frac{9}{5})\right)}\right|^{t}.$
In particular, making $t\to 1+,$ we get
$\sum_{f\in\mathcal{F}_{k}(N)}|\lambda_{n}(f)|\geq\frac{(4\pi)^{k-1}}{\Gamma(k-1)}\left|\frac{0.002397\pi}{(k-1)^{1/3}}e^{(k-1)\left(1-\frac{4}{9}-\log(\frac{9}{5})\right)}\right|.$
Although Corollary 2 is mentioned for a specific value of $D_{0}=0.999,$ and $\epsilon=0.4470$, the result holds for any $D_{0}\in(e^{H_{0}},1)$ and $0.44730<\epsilon<$ 0.4473 . In particular, Corollary 2 gives us a lower bound for the trace of the operator $(T_{n})^{t}$ for a positive integer $t>1,$ whenever $n$ satisfies the hypothesis of the corollary.
@@
Let $m,n,N$ be natural numbers with $(mn,N)=1$ and $N$ odd. Let $N=df$ where $d$ is powerful and $f$ is squarefree, where $(d,f)=1$. Let $\frac{4\pi\sqrt{mn}}{N}\in\left[(k-1),(k-1)+(k-1)^{\frac{1}{3}}\right]$ and assume that $mn$ is a quadratic residue modulo each prime that divides the powerful part of $N$. Let $E_{0}$ be a positive constant less than $0.327H(N)$ (as defined in Theorem 2) and $k_{1}(N):=2+\left\lfloor\frac{\log(1.632)-\log(0.327H(N)-E_{0})}{-H_{0}}\right\rfloor$. Then, for all $k\geq\max(180,k_{1}(N)),$ we have
@@
We obtain the following corollary to Theorem 3 which is an analogue of Corollary 2 with a sharper lower bound.
Corollary 3.
Let $k,\,N$ be natural numbers with $k\geq\max(22,k_{1}(N)),$ $k$ even and $N$ odd. Let $n\geq 2$ be a natural number such that $n\in\left[\left(\frac{(k-1)N}{4\pi}\right)^{2},\left(\frac{\left((k-1)+(k-1)^{\frac{1}{3}}\right)N}{4\pi}\right)^{2}\right]$ with $(n,N)=1$, and $n$ is a quadratic residue modulo each prime dividing the powerful part of $N$. Then for $t>1,$ we have
$\sum_{f\in\mathcal{F}_{k}(N)}|\lambda^{t}_{n}(f)|\geq\left(\frac{(4\pi)^{k-1}2\pi E_{0}}{\Gamma(k-1)}\right)^{t}\frac{(k-1)^{-\frac{t}{3}}}{|\mathcal{F}_{k}(N)|^{t-1}},$
where $E_{0}$ be a positive constant less than $0.327H(N)$. In particular, making $t\to 1+,$ we get
$\sum_{f\in\mathcal{F}_{k}(N)}|\lambda_{n}(f)|\geq\frac{(4\pi)^{k-1}2\pi E_{0}}{\Gamma(k-1)}\frac{1}{(k-1)^{1/3}}.$
Organization of the article
The article is organised as follows. In Section 2, we discuss the proof of the lower bound of classical Kloosterman sums. In Sections 3, we prove Theorems 2 and 3 and some corollaries. In the last section of the article, we provide a table for the values of $H(N),k_{0}(N)$ and $k_{1}(N)$ for some specific values of $N.$
@@ 2. Lower bound for Kloosterman sums
In this section, we primarily prove Theorem 1 by considering squarefree and odd prime power cases separately and we consider the general case in section 2.3.
@@ 2.1. Squarefree case
Let $\alpha=S(a,b;c)\in\mathbb{Q}(\zeta_{c})$, where $\zeta_{c}=e^{2\pi i/c}$. An element of Gal$(\mathbb{Q}(\zeta_{c})/\mathbb{Q})$ sends $\zeta_{c}$ to $\zeta_{c}^{n}$ where $(n,c)=1.$ Hence, the conjugates of $\alpha$ are the Kloosterman sums $S(an,bn;c),$ where $(n,c)=1.$ In fact, we have $S(an,bn;c)=S(an^{2},b;c)=S(a,bn^{2};c).$ For $c$ prime, $S(a,b;c)$ is a linear combination of at most $c-1$ $c$-th roots of unity, but they are necessarily linearly independent over $\mathbb{Q}$. So if the linear combination becomes 00, all the coefficients need to be 00 which is not the case. Hence $S(a,b;c)\neq 0$ for $c$ prime. Using the multiplicativity of Kloosterman sums (see Lemma 4), $S(a,b;c)\neq 0$ for squarefree $c$. Let $\mathcal{N}(\alpha)$ denote the norm of $\alpha\in\mathbb{Q}(\zeta_{c})$ over $\mathbb{Q}.$ Since $\alpha=S(a,b;c)$ is a non-zero algebraic integer in $\mathbb{Q}(\zeta_{c})$, its norm $\mathcal{N}(\alpha)$ is a non-zero integer and therefore,
@@ 2.2. Odd prime power case
Let $q$ be odd, and $\epsilon_{q}$ be $1$ or $i$ depending on whether $q\equiv$ $1$ or $-1$ modulo $4$ For odd prime power moduli $q$, we have the following explicit evaluation of $S(a,b;q)$.
@@
[IK04, Chapter 12, Exercise 1] Let $q=p^{\beta}$ for some prime $p$ and $\beta\geq 2.$ Let $p$ does not divide $2ab$. If $ab$ is a quadratic residue modulo $p$, then
@@
We start by noting that $q$ does not divide $l$ and hence $\left|\left(\dfrac{l}{q}\right)\right|=1.$ Let $l=qt+r$ where $10\leq r\leq q-$ 1 . Note that $p\nmid 2ab$ implies that $p$ is odd and so is $q.$ We consider the following two cases depending on whether $q\equiv$ $1$ or $-1$ modulo $4$
@@
We first show that $\left\|\frac{q-8r}{2q}\right\|\neq 0.$ Suppose not, then $\frac{q-8r}{2q}=t$ for some $t\in\{0,-1,-2,-3\}$ as $10\leq r\leq q-$ 1 . We have $p\,|\,r^{2}-ab$ which implies that $p$ divides $\frac{(1-2t)^{2}q^{2}-64ab}{64}.$ This is an impossibility since $p$ does not divide $2ab.$ A similar argument yields $\left\|\frac{4r}{q}\right\|\neq 0.$ Proceeding now from the equation (2.3), we conclude that
@@ 2.3. General Case
In this section, we apply the multiplicativity property of Kloosterman sums. For $m,n\in\mathbb{N},$ $(m,n)=1,$ let $\overline{m}(n)$ denote the multiplicative inverse of $m$ in $(\mathbb{Z}/n\mathbb{Z})^{*},$ i.e., $m\overline{m}(n)\equiv 1$ (mod $n$). We have the following factorization of Kloosterman sums into Kloosterman sums to prime power moduli.
@@
for $1\leq i\leq t.$ Further let $m_{0}=1,$ $m_{i}=\prod_{j=1}^{i}\overline{l_{j}}({c_{j}})$ for $11\leq i\leq t-$ 1 . Then
@@
We first factorize $S(a,b;c)$ using Proposition 4 into a Kloosterman sum with modulus $f$ and a product of Kloosterman sums with prime power moduli $l_{i}$ dividing $d$. Then we apply Proposition 1 to the Kloosterman sum with modulus $f$ and Proposition 3 to the Kloosterman sums with moduli $l_{i}$, taking the condition $(ab,c)=1$ and Remark 2 into account. ∎
@@
Let $a,b\in\mathbb{Z}$ and $c$ be an odd natural number such that $c=df,$ where $f$ is squarefree and $d$ is a powerful number with $(d,f)=1,$ and $(ab,c)=1$. Assume that $ab$ is a quadratic residue modulo each prime dividing $d$. Then
$\left|S(a,b;c)\right|\geq\left(\frac{f^{2}}{c}\right)^{1/2}\frac{2^{\omega(c)}}{(\tau(f)^{2}f)^{\phi(f)/2\tau(f)}}.$
@@
$\begin{split}&\frac{2^{\omega(d)}}{\sqrt{d}}\left(\frac{1}{\left(\tau(f)\sqrt{f}\right)^{\phi(f)/\tau(f)-1}}\right)\\ \geq&\frac{2^{\omega(d)+\omega(f)}}{\sqrt{df}}\left(\frac{1}{\left(\tau(f)\right)^{\phi(f)/\tau(f)}}\right)\left(\frac{1}{\left(\sqrt{f}\right)^{\phi(f)/\tau(f)-2}}\right)\\ \geq&\frac{2^{\omega(c)}}{\sqrt{c}}\frac{1}{\tau(f)^{\phi(f)/\tau(f)}}\frac{1}{f^{\phi(f)/2\tau(f)-1}},\end{split}$
where we have used $\tau(f)=2^{\omega(f)}$ for squarefree $f.$

### v4

@@ Abstract.
We present a lower bound for the classical Kloosterman sum $S(a,b;c)$ where $(ab,c)=1$ and $c$ is an odd integer. We apply this lower bound for Kloosterman sums to derive an explicit lower bound in Petersson’s trace formula, subject to a given condition. Consequently, we achieve a modified version of [JS20, Theorem 1.7], where weight $k$ and level $N$ are permitted to vary independently. Using this modified version, we get a lower bound for a weighted trace of the Hecke operator $T_{n}$ acting on the space $S_{k}(N)$, of cusp forms of weight $k$ and level $N$ with $(n,N)=1$.
@@ 1. Introduction
where $a,b$ are integers, $c$ is a natural number and $\overline{x}(c)$ denotes the multiplicative inverse of $x$ modulo $c$. The sums for the special case of $a=0$ or $b=0$ are called Ramanujan sums. Kloosterman sums appear in various problems in number theory. One of the important places of its occurrence is in the Kuznetsov trace formula [Kuz80, Theorem 1] and Petersson’s trace formula. Recently, the non-vanishing aspect of Kloosterman sums has been used in [BDS23, Theorem 2] and [Das24] to obtain discrepancy results for the distribution of eigenvalues of Hecke operators. Whenever we have non-vanishing of a sum, a natural question to ask is whether we can obtain a lower bound for the same. To the best of our knowledge, no such lower bound for Kloosterman sums has been established in the literature. In the next section, we answer this question for $S(a,b;c),$ where $(ab,c)=1$. First, we obtain the following lower bound for classical Kloosterman sums.
@@
Let $c$ be an odd natural number such that $c=du$, where $u$ is squarefree and $d$ is a powerful number with $(d,u)=1$. Let $a,b$ be integers such that $(ab,c)=1$ and $ab$ is a quadratic residue modulo each prime dividing $d$ (and hence, $ab$ is a square modulo $d$ by Hensel’s lemma and Chinese remainder theorem). Then
$\left|S(a,b;c)\right|\geq\frac{2^{\omega(d)}}{\sqrt{d}}\left(\frac{1}{\left(\tau(u)\sqrt{u}\right)^{\phi(u)/\tau(u)-1}}\right),$
where $\omega(d)$ is the number of distinct prime divisors of $d$, $\tau(u)$ is the number of positive divisors of $u,$ and $\phi(u)$ denotes the Euler’s phi function of $u$. If $ab$ is a quadratic non-residue modulo a prime dividing $d$ (and hence $ab$ is not a square modulo $d$), then $S(a,b;c)=0$.
@@
Let $m,n,N$ be natural numbers with $(mn,N)=1$ and $N$ odd. Let $H_{0}=1-\frac{4}{9}-\log\left(\frac{9}{5}\right)\approx-0.03223$, $D_{0}\in\left(e^{H_{0}},1\right)$ be a real number ($e^{H_{0}}\approx 0.96828)$. Let $N=du$ where $d$ is powerful and $u$ is squarefree, where $(d,u)=1$. Let $k_{0}(N)=2+\left\lfloor\frac{\log 7-\log H(N)}{\log A_{0}}\right\rfloor,$ where $A_{0}=D_{0}e^{-H_{0}},$ $H(N)=2^{\omega(N)}\frac{G(N)}{N^{3/2}}$ with $G(N)=\frac{u}{(\tau(u)^{2}u)^{\phi(u)/2\tau(u)}}.$ Let $k\geq\max(22,k_{0}(N))$, $\frac{4\pi\sqrt{mn}}{N}\in\left[D_{0}(k-1),(k-1)\right]$ and assume that $mn$ is a quadratic residue modulo each prime dividing $d$. Let $\Delta_{k,N}(m,n)$ be as defined in equation (3.1), then
@@
  • The theorems [JS20, Theorem 1.7] and [Das24, Theorem 4], assume $N$ to be fixed and provide an asymptotic for $|\Delta_{k,N}(m,n)-\delta(m,n)|,$ as $k\rightarrow\infty.$ However, Theorem 2 is applicable for every $k$ and $N$, provided they satisfy certain hypotheses. For a fixed level $N,$ [ILS00, Corollary 2.3] gives an asymptotic for $\Delta_{k,N}(m,n)$ with the condition $\frac{4\pi\sqrt{mn}}{N}\leq\frac{k}{3},$ whereas Theorem 2 gives us an asymptotic for the condition $D_{0}(k-1)\leq\frac{4\pi\sqrt{mn}}{N}\leq(k-1).$
@@
The following corollary is immediate.
@@
Let $0<\epsilon<0.4473.$ Let $m,n,N$ be natural numbers with $(mn,N)=1$ and $N$ odd. Let $H_{0},D_{0},k_{0}(N)$ be as defined in Theorem 2. Let $N=du$ where $d$ is powerful and $u$ is squarefree with $(d,u)=1$. If $\frac{4\pi\sqrt{mn}}{N}\in\left[D_{0}(k-1),(k-1)\right]$ and $mn$ is a quadratic residue modulo each prime dividing $d$, then for all $k\geq\max\left(k_{0}(N),1+\left(\frac{3}{\epsilon}\right)^{3/2}\right),$ we have
@@
Let $m,n,N$ be natural numbers with $(mn,N)=1$ and $N$ odd. Let $N=du$ where $d$ is powerful and $u$ is squarefree, where $(d,u)=1$. Let $\frac{4\pi\sqrt{mn}}{N}\in\left[(k-1),(k-1)+(k-1)^{\frac{1}{3}}\right]$ and assume that $mn$ is a quadratic residue modulo each prime that divides the powerful part of $N$. Let $E_{0}$ be a positive constant less than $0.327H(N)$ (as defined in Theorem 2) and $k_{1}(N):=2+\left\lfloor\frac{\log(1.632)-\log(0.327H(N)-E_{0})}{-H_{0}}\right\rfloor$. Then, for all $k\geq\max(180,k_{1}(N)),$ we have
@@
Organisation of the article
The article is organised as follows. In Section 2, we discuss proofs for the lower bound of classical Kloosterman sums and for Corollary  2. In Sections 3, we prove Theorems 2, 3 and Corollary  3. In the very end of Section 3, we provide a table for the values of $H(N),k_{0}(N)$ and $k_{1}(N)$ for some specific values of $N.$
@@ 2. Lower bound for Kloosterman sums
In this section, we primarily prove Theorem 1 by considering squarefree and odd prime power cases separately, and we consider the general case in Section 2.3.
@@ 2.1. Squarefree case
[…]
@@ 2.2. Odd prime power case
Let $q$ be odd, and $\epsilon_{q}$ be $1$ or $i$ depending on whether $q\equiv$ 11 or 1-1 modulo $4.$ For odd prime power moduli $q$, we have the following explicit evaluation of $S(a,b;q)$.
@@
[IK04, Chapter 12, Exercise 1] Let $q=p^{\beta}$ for some prime $p$ and $\beta\geq 2.$ Suppose $p$ does not divide $2ab$. If $ab$ is a quadratic residue modulo $p$, then
@@
We start by noting that $q$ does not divide $l$ and hence $\left|\left(\dfrac{l}{q}\right)\right|=1.$ Let $l=qt+r$ where $0\leq r\leq q-1.$ Note that $p\nmid 2ab$ implies that $p$ is odd and so is $q.$ We consider the following two cases depending on whether $q\equiv$ 11 or 1-1 modulo $4.$
@@
We first show that $\left\|\frac{q-8r}{2q}\right\|\neq 0.$ Suppose not, then $\frac{q-8r}{2q}=t$ for some $t\in\{0,-1,-2,-3\}$ as $0\leq r\leq q-1.$ We have $p\,|\,r^{2}-ab$ which implies that $p$ divides $\frac{(1-2t)^{2}q^{2}-64ab}{64}.$ This is an impossibility since $p$ does not divide $2ab.$ A similar argument yields $\left\|\frac{4r}{q}\right\|\neq 0.$ Proceeding now from the equation (2.3), we conclude that
@@ 2.3. General Case
In this section, we apply the multiplicativity property of Kloosterman sums. For $m,n\in\mathbb{N},$ $(m,n)=1,$ let $\overline{m}(n)$ denote the multiplicative inverse of $m$ in $(\mathbb{Z}/n\mathbb{Z})^{*},$ that is, $m\overline{m}(n)\equiv 1$ (mod $n$). We have the following factorization of Kloosterman sums into Kloosterman sums to prime power moduli.
@@
for $1\leq i\leq t.$ Further let $m_{0}=1,$ $m_{i}=\prod_{j=1}^{i}\overline{l_{j}}({c_{j}})$ for $1\leq i\leq t-1.$ Then
@@
We first factorise $S(a,b;c)$ using Proposition 4 into a Kloosterman sum with modulus $u$ and a product of Kloosterman sums with prime power moduli $l_{i}$ dividing $d$. Then we apply Proposition 1 to the Kloosterman sum with modulus $u$ and Proposition 3 to the Kloosterman sums with moduli $l_{i}$, taking the condition $(ab,c)=1$ and Remark 2 into account. ∎
@@
Let $a,b\in\mathbb{Z}$ and $c$ be an odd natural number such that $c=du,$ where $u$ is squarefree and $d$ is a powerful number with $(d,u)=1,$ and $(ab,c)=1$. Assume that $ab$ is a quadratic residue modulo each prime dividing $d$. Then
$\left|S(a,b;c)\right|\geq\left(\frac{u^{2}}{c}\right)^{1/2}\frac{2^{\omega(c)}}{(\tau(u)^{2}u)^{\phi(u)/2\tau(u)}}.$
@@
$\begin{split}&\frac{2^{\omega(d)}}{\sqrt{d}}\left(\frac{1}{\left(\tau(u)\sqrt{u}\right)^{\phi(u)/\tau(u)-1}}\right)\\ \geq&\frac{2^{\omega(d)+\omega(u)}}{\sqrt{du}}\left(\frac{1}{\left(\tau(u)\right)^{\phi(u)/\tau(u)}}\right)\left(\frac{1}{\left(\sqrt{u}\right)^{\phi(u)/\tau(u)-2}}\right)\\ \geq&\frac{2^{\omega(c)}}{\sqrt{c}}\frac{1}{\tau(u)^{\phi(u)/\tau(u)}}\frac{1}{u^{\phi(u)/2\tau(u)-1}},\end{split}$
where we have used $\tau(u)=2^{\omega(u)}$ for squarefree $u.$